if the non zero vectors a and b are perpendicular to each other,then the solution of the equation r x a=b willbe....

Grade:11

2 Answers

ng29
9 years ago
plz check the statement again and post it again
Lôvë Àññy
6 years ago
[r = xa +yb + z(a×b)]From given equation we know r perpendicular to b thus y= 0,>[a×r = a×xa + a×z(a×b),]>(-b = z[(a.b)a - |a|²b])>(z=1/|a|²)>(r= xa +( a×b)/|a|²)